The ratio of longest wavelength and the shortest wavelength observed in the five spectral lines of emission spectrum of hydrogen is 
a)4/3
b)525/376
c)25
d)900/11

Grade:Select Grade

4 Answers

Anjali Chowdary
9 years ago
The shortest one is from the transition from n=11 at 91.9nm
the longest transitionis from a transition from n=2 at 121.6nm
Using Rydberg formula.
1/lambda=R*(1-1/n^2),where n=2,3,4,....
In that case the largest value of 1/lambda(smallest value of lambda)would be the smallest value of n.
 
Karan Shinde
9 years ago
you are correct. Please help me with the solution of the problem...
Karan Shinde
9 years ago
the answer is d)
Casintra Fernandes
6 years ago
For longest wavelength, ni=1 & nf=2.
Using 1/lambda=R [1/(ni)^2 – 1/(nf)^2]
substituting the above you will get, lambda1 =4/3R
For shortest wavelength, ni=1 & nf=infinity.
Using 1/lambda=R [1/(ni)^2 – 1/(nf)^2]
substituting the above you will get, lambda2 =1/R
Therefore lambda1/lambda2 =4/3
So answer should be (a)